Circular No. 6733

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discovery of two more probable Kreutz sungrazing comets. Measurements by St. Cyr and D. Biesecker have been reduced by G. V. Williams and are given on MPEC 1997-R03 and 1997-R04, together with tentative orbital elements by B. G. Marsden. C/1997 Q1, with no evidence of a tail and peak magnitude ∼ 7, was discovered by Biesecker in C3 coronagraph data, which were interrupted when the comet was at an apparent solar distance of ∼ 12.5 R_{⊙}, 13 hr after the first observation, due to the closing of the telescope doors. C/1997 Q2, discovered by K. Schenk in C3 data nine days earlier, is one of the brighter SOHO discoveries, peaking at mag 4 or brighter; it developed a tail, which was at least 1^{o}.5 long in the C2 images, 26–30 hr after the initial C3 detection. As with other SOHO comets with extensive observed arcs, the orbital solution for C/1997 Q2 does not fit the complete arc, but Marsden notes that the two comets were clearly in the same Kreutz-type track, with T = Sept. 1.3 and Aug. 24.1 TT, respectively.

1996 UT  α_{2000}  δ_{2000}  Comet
Aug. 31.399  10^{h}15.^{m}1  + 7^{o}16′  C/1997 Q1
22.771  9 39.3  + 9 52  C/1997 Q2

GRB 970828 A. Guarnieri, C. Bartolini, N. Masetti, and A. Piccioni, University

of Bologna, communicate: “A preliminary analysis of CCD images of the GRB 970828 field (IAUC 6728), obtained with the Bologna University 1.5m telescope (+ RCA) in B and R bands (limiting mag ∼ 21 for both) on Aug. 29.86, 30.84, and 31.83 UT, and in the R band on Sept. 1.82, shows no object with a rise or a decline of > 0.2 mag.”

SUPERNOVA 1997dc IN NGC 7678 C. Y. Peng, University of Arizona; R. R. Treffers, M. Modjaz, and A. V.

Filippenko, University of California at Berkeley; and M. W. Richmond, Rochester Institute of Technology, report that SN 1997dc (cf. IAUC 6715– 6717) is visible in prediscovery CCD images obtained as part of the Lick Observatory Supernova Search (cf. IAUC 6627) on July 26 and Aug. 1 and 5 UT. It might be present on July 21, as well. The object was in all cases fainter than the threshold for automatic detection.

1997 September 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green